OpenStudy (anonymous):

Solve –3x2 + 30x – 90 = 0 please help i give medals! Answer choices: A.x=5+/-2i√5 B.x=5+/-i√5 C.x=10+/-2i√5 D.x=10+/-i√5 First answer gets medal

OpenStudy (anonymous):

PLease help

OpenStudy (anonymous):

Quadratic formula

OpenStudy (anonymous):

i know that lol im not sure how to actually work it

OpenStudy (anonymous):

cwrw 238 are you writing anything?

OpenStudy (cwrw238):

divide through by -3 x^2 - 10x + 30 = 0 using the quadratic formula: x = [-(-10) +/1 (sqrt(100 - 4*1*30)] / 2 = 10 +/- (sqrt( -20) / 2 = [10+/- sqrt(4 *-5)] / 2 = 5 +/- i√5 where i = sqrt(-1)

OpenStudy (anonymous):

Thanks!